OpenStudy (anonymous):

Find the finite number which (sqrt2/2)^n-1 approaches. Please include explanation.

OpenStudy (asnaseer):

is this the question?\[\lim_{n\rightarrow\infty}{(\frac{\sqrt{2}}{2})^n-1}\]

OpenStudy (anonymous):

I don't know what that is...

OpenStudy (anonymous):

what is bigger, \(\sqrt{2}\) or \(2\)?

OpenStudy (anonymous):

2 is bigger.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

because if you have a faction that is less than one and you raise it to higher and higher powers it gets smaller and smaller

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so since \(\frac{\sqrt{2}}{2}<1\) we know \[\lim_{x\to \infty}(\frac{\sqrt{2}}{2})^n=0\]

OpenStudy (anonymous):

guess i meant \[\lim_{n\to \infty}(\frac{\sqrt{2}}{2})^n=0\]

OpenStudy (anonymous):

wats the lim thing?

OpenStudy (anonymous):

and why does that equal zero?

OpenStudy (anonymous):

If I'm taking powers of a, that is, a^n there are basically three things that can happen. The simplest case is when a=1 if a = 1, then 1^1 = 1 1^2 = 1 1^3 = 1 and so on. Any power of a will just give me 1. The second case is that a is bigger than 1. For example, let's see what happens when a is 2 2^1 =2 2^2 = 4 2^3 = 8 2^4 = 16 So, in general if a is bigger than 1, then powers of a will get bigger and bigger, and they'll grow faster and faster as we take higher powers of n. The third case is that a is less than 1. For example, let's see what happens when a is (1/2) (1/2)^1 = (1/2) (1/2)^2 = (1/4) (1/2)^3 = (1/8) (1/2)^4 = (1/16) So, in general, if a is less than 1, then powers of a will get smaller and smaller as we take higher powers. And they get really small really quickly, giving us tiny fractions that are almost 0.

OpenStudy (anonymous):

you are asking what it approaches as n gets large (i assume) the succinct way to write that in math is \(\lim_{n\to \infty}\)

OpenStudy (anonymous):

Here is for example a the sequence \[ \left\{n,\left(\frac{2}{3}\right)^n\right\} \] for n =1 to 20 \[\left( \begin{array}{cc} 1 & 0.666667 \\ 2 & 0.444444 \\ 3 & 0.296296 \\ 4 & 0.197531 \\ 5 & 0.131687 \\ 6 & 0.0877915 \\ 7 & 0.0585277 \\ 8 & 0.0390184 \\ 9 & 0.0260123 \\ 10 & 0.0173415 \\ 11 & 0.011561 \\ 12 & 0.00770735 \\ 13 & 0.00513823 \\ 14 & 0.00342549 \\ 15 & 0.00228366 \\ 16 & 0.00152244 \\ 17 & 0.00101496 \\ 18 & 0.000676639 \\ 19 & 0.000451093 \\ 20 & 0.000300729 \\ \end{array} \right) What do you think that limit is hwn n goes to infinity? \]

OpenStudy (anonymous):

What do you think that limit is when n goes to infinity?

OpenStudy (anonymous):

\[ \left(\frac{2}{3}\right)^{28 }=0.000011734 \]

OpenStudy (anonymous):

oh... But the real questoin is does the sum of the perimeters approach a finite number? If so, what number? And the equation that I found for the perimeters is (sqrt 2/2)^n-1. So the first square that I am taking the perimeter of is 4 in perimeter.

OpenStudy (anonymous):

This happens as it is explained above by @SmoothMath because 2/3 < 1

OpenStudy (anonymous):

The first square has a length of 1 and there is another square inside of it with its corners on the midpoint of each side of the first square. and it goes on and on.

OpenStudy (anonymous):

Your formula is not accurate since the limit of your formula is negtive and the perimeter is positive.

OpenStudy (anonymous):

How do you know that? How do you find the limit?

OpenStudy (anonymous):

the actual equation I should've gotten is 4x(sqrt 2/2)^n-1), isn't it?

OpenStudy (anonymous):

Let us call the perimeters \[p_0 , p_1, .... p_n... \] It is easy to show that the second perimeter \[ p_1= \frac {p_0}{\sqrt 2} \] and in general \[ p_n= \frac {p_{n-1}}{\sqrt 2} \] Do we agree on that?

OpenStudy (anonymous):

what??? I have no idea what you typed... uh

OpenStudy (anonymous):

So \[ p_n = \frac {p_0}{(\sqrt 2)^n}= \frac {1}{(\sqrt 2)^n} 4= 4 r^n\\ \text { where } \\ r= \frac 1 {\sqrt 2} \] The sum is \[ 4\sum_{n=0}^\infty r^n \] This is a geometric series. Do you know how to find its sum? Notice r<1

OpenStudy (anonymous):

Every new perimeter is obtained from the previous one by dividing the previous perimeter by \[ \sqrt 2\]

OpenStudy (anonymous):

how do you translate this? do you put it into the equation thing with the sigma right under the comment box?

OpenStudy (anonymous):

Have you heard about geometric series?

OpenStudy (anonymous):

Read this http://en.wikipedia.org/wiki/Geometric_series

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what is a + ar + ar^2 + .... + a r^n ?

OpenStudy (anonymous):

a(r)^n-1

OpenStudy (anonymous):

I just don't know how to find the limit

OpenStudy (anonymous):

It is \[ \frac{a \left(1-r^{n}\right)}{1-r } \]

OpenStudy (anonymous):

I dont know what\[ \frac{a \left(1-r^{n}\right)}{1-r } \] is because I can't see what it actually klooks like and I can't make sense of it... I'm sorry

OpenStudy (anonymous):

If r < 1, then the limit when n goes to infinity of \[ r^n \] is zero

OpenStudy (anonymous):

Read the link that I gave you before to understand what is going on?

OpenStudy (anonymous):

But it says here that the answer is 8 +4 sqrt2

OpenStudy (anonymous):

Yes it is true the answer is \[ 8+4 \sqrt{2} \]

OpenStudy (anonymous):

how...

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Because the limit when n goes to infinity of \[ \frac{a \left(1-r^{n}\right)}{1-r } \] is \[ \frac{a \left(1-0\right)}{1-r }= \frac 4{ 1 - \frac 1 {\sqrt 2}}=8+4 \sqrt{2} \]

OpenStudy (anonymous):

After simplifying

OpenStudy (anonymous):

My friend just told me to use conjugates. Could you explain conjugates to me?

OpenStudy (anonymous):

wow this turned out to be a totally different question didn't it?

OpenStudy (anonymous):

\[ \frac{4}{1-\frac{1}{\sqrt{2}}} =\frac{4}{1-\frac{\sqrt{2}} {2}}=\frac{4}{\frac{1}{2} \left(2-\sqrt{2}\right)}=\frac{8}{2-\sqrt{2}} \]

OpenStudy (anonymous):

Multiply up and down by the conjugate \[ 2 + \sqrt 2 \] You obtain \[ \frac{8}{2-\sqrt{2}} \frac { 2+ \sqrt 2}{2+ \sqrt 2}= \frac {8( 2 + \sqrt 2)} {4-2}= 4 (2 +\sqrt 2)= 8 + 4 \sqrt 2 \]

OpenStudy (anonymous):

Thank you for all of your time,. I think I get how to do it now!!!